### v1

# A Prompt-Engineered Large Language Model, Deep Learning Workflow for Materials Classification

Abstract
With the advent of ChatGPT, large language models (LLMs) have demonstrated considerable progress across a wide array of domains. Owing to the extensive number of parameters and training data in LLMs, these models inherently encompass an expansive and comprehensive materials knowledge database, far exceeding the capabilities of individual researcher. Nonetheless, devising methods to harness the knowledge embedded within LLMs for the design and discovery of novel materials remains a formidable challenge. In this study, we introduce a general approach for addressing materials classification problems, which incorporates LLMs, prompt engineering, and deep learning algorithms. Utilizing a dataset of metallic glasses as a case study, our methodology achieved an improvement of up to 463% in prediction accuracy compared to conventional classification models. These findings underscore the potential of leveraging textual knowledge generated by LLMs for materials especially with sparse datasets, thereby promoting innovation in materials discovery and design.
The concept of “AI for Materials” refers to the utilization of artificial intelligence (AI) techniques, such as machine learning (ML), deep learning (DL), and the increasingly prevalent large language models (LLMs), to design novel materials or investigate the composition-structure-property relationships in different materials [1, 2, 3, 4]. Significant strides have been achieved in fields such as biomaterials [5, 6] and organic materials [7, 8], which can be attributed to the relatively uniform representation of organic molecules [9] and the availability of numerous high-quality datasets [10]. However, for inorganic materials, particularly those with large compositional space (element types 4), the application of AI presents a more complex challenge. This complexity arises due to factors such as the scarcity of experimental data [4], diverse properties of inorganic materials, complicated crystal structures, potential existence of intermediate phases (e.g. intermetallics in alloys), and defects including vacancies, dislocations, and grain boundaries [11, 12]. Contrasting with organic molecules that can be modeled with neural networks, multi-component inorganic materials often involve the amalgamation of various phases and defects or even the formation of new phases, presenting an ongoing challenge in the structural representation [13, 14]. Although large-scale databases such as the Materials Project [15], ICSD [16], and AFLOW [17] exist, data pertinent to specific tasks such as the relationships between composition and mechanical properties in alloys and structural stability of two-dimensional materials are often sparse and dispersed across different datasets, complicating the assembly of substantial databases for training purposes. Furthermore, input features can vary across datasets, sometimes necessitating manual feature construction based on domain knowledge [18]. Collectively, these issues hinder the application of AI in the realm of inorganic materials, subsequently impeding the discovery and design of new materials.
The advancement of LLMs, particularly with the advent of ChatGPT, has sparked a surge of interest in distilling knowledge from these models through prompt engineering [19, 20, 21]. Notable examples include constructing scientific question answering knowledge bases using LLMs [22], transferring knowledge from large to small models via in-context learning [23], and training materials-specific LLMs [24, 25]. Given the expanding capabilities of LLMs and the growing volume of trained data, it is conceivable to consider these models as encyclopedia resources for materials science, enabling the extraction of text-based knowledge [26]. As shown in the upper part of Fig. 1, this approach addresses some challenges associated with data collection and feature extraction in conventional processes. By representing all content in textual format, the generated data becomes universally applicable, offering a versatile solution for materials science research.
Therefore, we have established a universal workflow for material text feature-label classification using textual data generated by LLMs, as illustrated in the lower part of Fig. 1. The workflow is structured as follows: (i) define the material classification problem to be addressed; (ii) design prompts via prompt engineering to distill knowledge from LLMs and store it as textual data; (iii) fine-tune a bidirectional encoder representations transformers (BERT) [27] model (commonly employed in natural language processing) to train on the textual data-label pairs; (iv) apply the model to explore new materials or study composition-structure-property relationships. In an example problem involving the classification of 5,577 metallic glasses (MGs) with experimental datasets labeling the MG forming categories (bulk MG, ribbon, or non-ribbon), a BERT model trained with optimized prompts and position embedding methods achieved up to a 48% increase in classification accuracy compared to traditional ML models. The classification models obtained from our workflow achieved an overall accuracy of 97.7%. For the classification of bulk MGs (BMGs) with the smallest sample number (11% of the entire dataset), the accuracy improved by up to 463% compared to traditional ML models. These results underscore the superiority of our workflow in addressing material classification problems and this workflow can be extended to various material applications. Additionally, tremendous potential of “AI for Materials” is revealed, grounded in natural language processing and prompt engineering.
Text-based materials
classification workflow
Our material classification workflow comprises the four-tier structure depicted within the blue box in Fig. 1, and is readily adaptable to various material classification problems. Initially, researchers are required to undertake fundamental data collection and delineate the prediction objectives and categories, such as determining whether a material exhibits stability, corrosion resistance, or favorable optoelectronic transmission properties.
Secondly, the inputs for ML models in conventional workflows typically comprise material properties obtained through experiments or theoretical calculations, demanding substantial expertise as well as experimental or simulation efforts. Different researchers may select varying property features, leading to multiple datasets for the same problem [28]. However, training numerical-based ML models on the combination of different datasets often fails due to inconsistencies in input feature vectors. Furthermore, the number of datasets for specific niche tasks in materials science research is quite small ( several hundreds), negligible compared to the millions of samples in classical computer science fields like ImageNet dataset [29]. For example, in the field of alloys, the dataset for fracture and impact toughness of high-entropy alloys contains only 154 data points [30], the fatigue database of complex metallic alloys consists of merely 272 data points [31], and the dataset of mechanical properties of high-entropy alloys has 370 data points [32]. These limitations hinder the application of ML in alloys. In our workflow, we employ textual inputs generated by LLMs, offering the advantage of eliminating the need for manual feature extraction. Moreover, LLMs can output content relevant to specific problems in materials science with customized prompts. As LLMs possess considerably more parameters than traditional ML models, they can generate a richer and more diversified knowledge base.
Furthermore, within our workflow, the classification BERT model employs the attention mechanism for self-instruct learning, enabling automatic feature extraction. Different pre-trained models can be utilized for corresponding downstream tasks. Models based on the attention mechanism are not limited to text input but also applicable to areas like image recognition [33] and molecular structure identification [34]. These features allow our workflow to seamlessly support the combination of different datasets and multimodal expansion, addressing the issue of insufficient datasets for specific material problems often encountered in traditional methods.
Finally, the material classification models trained through our workflow can be utilized in numerous downstream tasks, such as discovering new materials across different categories, conducting interpretability analysis to extract composition-structure-property relationships of materials, and even fine-tuning models to achieve excellent results in other classification tasks. Our workflow represents a novel and general approach for material sciences by constructing large-scale databases from prompt-engineered LLMs and training general DL models to expedite materials discovery. We showcase the efficiency and accuracy of the workflow by using the classification problem of MGs as an example.
Application: metallic glasses classification
MGs are amorphous materials typically composed of a combination of metal and other elements [35]. They are renowned for superior engineering performance, including high strength, high thermal stability, and excellent corrosion resistance [36]. Despite ongoing efforts to discover MGs with enhanced comprehensive performance, their formability remains a complex and challenging problem in materials science. Predicting the glass formability is also difficult for atomistic simulations, due to the large size and long timescale (unattainable by first-principles calculations) and the absence of accurate interatomic potentials for molecular dynamics simulations. Moreover, MGs have an extensive compositional space, often comprising three or more elements, rendering the experimental search for new applicable MGs a time-consuming process [37, 38].
Various ML methods [39] have been employed to classify MGs; however, these methods often rely on manually calculated or carefully designed material features. The inconsistency of designed features and datasets across different ML models necessitates researchers to recalculate those features each time when utilizing other models. In the worst-case scenario, certain features rely on experimental calculations, rendering a MG dataset applicable to one model but not another. In this regard, we leveraged the experimental dataset of MGs by [40] and applied it to the proposed workflow here. The dataset compiled by [40] consists of 8,415 alloy compositions with their corresponding MG forming categories. The alloy compositions were classified into three categories based on the critical cooling rate (): BMG (), ribbon (), and non-ribbon (NR, ). The glass formability is then BMG ribbon NR. From this dataset, 5,577 alloy compositions with only a single glass-forming category were selected as input dataset, with 80% of them being train set while the rest being test set. As shown in Fig. 2a, a customized version of the general workflow was developed for the given input data. The blue arrows indicate the training process, which involves expanding the alloy compositions into textual data using an LLM. Subsequently, a DL model is trained on the textual data from the training set for MG classification. Finally, the analysis of efficiency and accuracy and model interpretation are performed using textual data from the test set. These three main steps are illustrated in Fig. 2b. The green arrows indicate the application stages of the model. For an unexplored alloy composition, the workflow can determine the MG category (BMG, ribbon, or NR) within seconds.
[…]
Next, we trained a MGs BERT (MgBERT) classification model that combines self-instruct learning and supervised classification. DL models for natural language processing typically require a vast number of fitting parameters, necessitating the selection of a suitable pre-trained model as the base model. MgBERT is fine-tuned based on MatSciBERT (a pre-trained model based on materials science texts) [42]. Figs. 4a, b display a comparison of different pre-trained BERT models, with MatSciBERT achieving the best results in classification accuracy and model loss on the test set. However, MatSciBERT can only accept inputs of up to 512 tokens, while our text data has a maximum length of 859 when converted to tokens. Consequently, we designed a dynamically resizable embedding layer called “MgBERT Embedding” to meet the requirements of input scalability. This embedding layer can assign weights based on layers and proportions according to the preset length of the new token, building upon the embedding weights trained by MatSciBERT. This approach generates new position embedding that can support longer text inputs and circumvent time-consuming weight initialization training. This new position embedding layer (in Fig. 4c) is utilized for tokenizing the input text of the model. From the comparison of results in Figs. 4g and h, our MgBERT embedding Method 3 (Fig. 4f) exhibits higher accuracy and lower train set loss, compared to the embedding layer method in Fig. 4d, which uses newly initialized weights, and that in Fig. 4e, which only employs pre-trained weights to cover the first 512 embedded positions. For different pre-trained models trained on the same text data (Figs. 4i, j), our MgBERT has the best performance in training set loss, training set classification accuracy, and test set classification accuracy. Notably, for alloy data in the test set that our model has never encountered before, our model achieved an accuracy of 88.5%, an improvement of 9% compared to Longformer, and an improvement of 4% compared to the best performing MatSciBERT. The detailed classification results can be observed in the confusion matrix in Supplementary Fig. S6. These findings showcase the success of our improvement in embedding structure, as it simultaneously supports longer input text and enhances accuracy. To examine the effectiveness of the entire workflow, we trained a baseline model based on logistic regression (LR) using the training set, as well as a support vector machine (SVM) model and a gradient boosting decision tree (GBDT) model. Then, we conducted tests on the test set and compared the classification performance of different methods (Fig. 5). For the default three-class problem, our workflow achieved a maximum accuracy increase of 48% across the entire dataset and a peak enhancement of 32% on the test set. Even more astonishing is our performance with respect to identifying BMGs, as shown in Fig. 5b. For this task, our accuracy soared by up to 463% on the entire dataset and by as much as 307% on the test set. Considering that BMGs constitute only 11% of our dataset, these results highlight the remarkable capability of our text-based material classification workflow in accurately recognizing patterns in data-scarce scenarios, surpassing traditional ML models.
DL models are often considered to lack algorithmic transparency [43]. Language models such as BERT typically have hundreds of millions of parameters [27], making it challenging to provide detailed explanations of the model at the algorithmic level. Here, we utilized the local interpretable model agnostic interpretations (LIME) [44] method to analyze from an input-output perspective. We selected text data of alloy compositions CuZrGa (classified as BMG), AgAlLa (classified as ribbon), and AlMnSi (classified as NR) as three examples. Fig. 6a and Supplementary Fig. S7 illustrate that alloy composition is the primary contributing factor to these classification results, consistent with the classical Inoue rule [45] and high-throughput experiment results [46]. According to Inoue’s rule, components of an alloy introduce varying atomic size differences, altering the internal dense random packing structure. Moreover, the model has extracted some elemental property information. For example, “Melting” refers to the melting point, which is one of the critical criteria for determining the formation of MGs [47]. Additionally, “745” is the ionization energy of copper, which could affect the interatomic forces within the alloy and thereby influence MG formation. Furthermore, “2477” represents the boiling point of gallium, and the polymer-like thermoplastic behavior observed in MGs is thought to be related to boiling points [48], also connected to the formation of functional MGs. These findings indicate that our text classification model can perform self-instruct learning and automatically extract features from a given text. Building on this foundation, in contrast to traditional ML models that rely on manually designed features as input, our text-based workflow enables models to accept raw text as input. This flexibility allows for rapid transfer and recycling of pre-trained model parameters between different materials classification tasks. We also visualized the attention scores within the classification model and the model extracts features based on three distinct patterns (Fig. 6b). This process ensures that after 12 layers of training, the output [CLS] (classify) token possesses sufficient information for classification. Since we use the value of the [CLS] token from the final layer to make judgments in the classification layer, Fig. 6c visualizes the emphasis placed on different sections of the text. The results suggest that the composition and alloy layers have garnered significant attention. A possible explanation is that key tokens within these two layers have captured essential features from the input text. Consequently, the [CLS] token can achieve accurate classification results by focusing on those specific token positions. These findings demonstrate that the classification MgBERT model, generated by our general workflow, has the abilities of dynamic feature extraction, modeling inter-word relationships, hierarchical representation learning, and bidirectional context understanding, which may be potential indications for good classification performance.
Discussion
We emphasize that classification cannot be successfully achieved using direct questioning alone (as seen in Supplementary Figs. S8-S10). Therefore, as part of our workflow, we designed a BERT-based classification model to categorize materials using text generated by the LLM. In practice, employing a classification model is a necessary step at this juncture, as directly querying the LLM yielded a 0 accuracy rate.
The general workflow here offers a novel approach to address a significant challenge in “AI for material science”, where datasets are often small. By utilizing a pre-trained BERT model, such as MatSciBERT optimized for material science, we can fine-tune it (pre-trained model) with textual data from prompt-engineered LLM to achieve high accuracy in classification problem (up to 99.0% for the entire dataset).
We propose a general approach for materials classification and demonstrate its efficiency in an example case of the classification problem in MGs. Our general workflow can be applied to any type of material applications, provided that labeled material samples are available. The findings point towards a future where the integration of advanced language models and domain-specific fine-tuning could revolutionize the material classification process, especially in cases with sparse data. We envision that LLMs may eventually serve as “world models” [49], encapsulating knowledge across diverse materials and domains. Consequently, such models based on text for classification, prediction, or even generating new materials will become potential tools for addressing challenges in material design. A sufficiently robust LLM could potentially provide desired classification outcomes directly through appropriate prompting and significantly streamline the process of materials discovery and design.
Methods
Data processing for metallic glasses dataset
In processing the dataset for metallic glasses, we curated our initial compilation of 8,415 entries by excluding alloys comprising multiple categories. This refinement resulted in a dataset encompassing 5,577 single-class alloys, along with their respective classifications. The aim was to construct an unambiguous dataset, thereby facilitating the accuracy of analysis and model training.
To implement prompt engineering, we harnessed the capabilities of the langchain library. We designed prompts with varying templates in advance, establishing a “Composition” variable to dynamically cycle through the list of alloy compositions. This approach enabled us to systematically generate textual data in a large scale, ensuring consistency and efficiency in our data preparation workflow.
For the ingestion of data into the MgBERT model, we divided the 5,577 data points into a training and test set at a ratio of 80:20. This arrangement yielded 4,460 entries for training constituting 80% of the dataset, and 1,117 entries for testing purposes. It was imperative that the MgBERT model was trained comprehensively, indicating that the data allocation is reasonable.
Regarding the evaluation of datasets, while the MgBERT model was assessed using the designated test set, the machine learning models underwent evaluation using a different approach because they can not receive textual data. We employed the Matminer and Pymatgen tools to convert the chemical formulas of alloys into “composition features” as defined within Matminer’s framework. However, it is noteworthy that three entries failed to transform correctly. Consequently, the test set for evaluation purposes comprised of 1,116 entries, reducing the overall dataset to 5,574 records. Compared to MgBERT, there were three fewer data sets, which had little impact on accuracy comparison.
Large language model and prompt engineering
In this work, we used the Google Gemini-pro large language model, which is currently free and surpasses human experts in massive multi-task language understanding [50]. As discussed earlier, our research harnesses the automated orchestration capabilities of the langchain library to script the bulk generation of prompts and textual data. This scripting forms an integral part of a generic workflow and offers a high degree of adaptability. The prompt templates we have developed are not rigid constructs; rather, they are designed to be inherently flexible, allowing for arbitrary modifications tailored to the specific demands and objectives of researchers. This level of customization is critical in the context of materials science where the nuances of the subject matter can vary significantly from one study to another. The ability to fine-tune prompts to align with particular research questions or datasets ensures that the language model can be effectively leveraged to generate effective and contextually relevant textual data outputs.
Our investigation sought to demonstrate the effectiveness of prompt engineering by creating a range of alternative prompt templates to act as comparative benchmarks. These templates comprised one that integrated a chain-of-thoughts (CoT) approach with a few-shot learning paradigm (denoted as “few-shot with CoT”), another that depended exclusively on few-shot learning examples (denoted as “few-shot”), a third that was based on zero-shot learning scenarios (denoted as “direct inquiry”), and our custom-designed “MetalPrompt” from our established workflow. By leveraging these diverse templates, we crafted specific prompts that were subsequently employed to interrogate the Gemini-pro model. The underlying rationale for this varied prompt strategy was to gather and juxtapose the accuracy metrics derived from different prompting methodologies.
MgBERT training and evaluation
Here we used MatSciBERT as the basic model and implemented a subsequent classifier function with MgBERT embedding to build MgBERT. Accuracy and cross entropy are used to evaluate the effectiveness of the model. For the classification principle of MgBERT, we used the value of [CLS] token in the last layer of the attention section of MgBERT as the output to aggregate the information of the entire model. Then, through a classifier, we compress the value of the token into a vector of length 3, and label 0 as BMG, 1 as Ribbon, and 2 as Non-ribbon. Then we calculate the maximum position of the output vector value and locate it to the final classification result. The overall expenses associated with developing MgBERT comprise $46 USD for training on Nvidia V100 (30 hours and $1.53 USD per hour) and $1.5 USD for 1.89 million input and 2.14 million output tokens (according to the official pricing of Gemini-pro on January 29, 2024). Upon applying MgBERT for inference on 1,000 different compositions, the cost is as low as $0.3 USD. This cost-effective nature of training and implementing MgBERT enables researchers to adopt this general approach across various materials.
Interpretability method
To elucidate the relationship between input and output, we used the LIME (local interpretable model agnostic explanations) technique, which is widely used to explain the predictions of machine learning models by constructing the relationship function via interpolation sampling. Here we displayed the top 25 important features and performed 1,000 interpolation samples for function fitting. Although our parameters were initially set to select the top 50 features, due to visualization constraints, we chose the top 25 features for display.
For the explanation at the attention level, we used the bertviz library to complete the attention score flow diagram for 12 layers and 12 attention heads. To attribute the importance of the [CLS] token in the final attention layer, we first averaged the attention score of the 12 attention heads in the last layer. Then, we divided the input text words into three layers based on their positions, namely, composition, element, and alloy layer. To evaluate the importance of each layer, we calculated the sum of attention scores within that layer, weighted by the token length ratio, and then divided by the total token length.
Data availability
The method section provides the models and algorithms employed in this study, while specific parameter implementations can be found in the supplementary notes in supplementary information. The data used in this work is available in the supplementary data, including the metallic glasses dataset, alloy description files generated by large language model Gemini-pro, MgBERT classification model training logs, and data used for figure plotting.
Code availability
All the codes used in the paper will be made available on GitHub upon the acceptance of manuscript, and these codes can also be acquired upon reasonable requests.
ReferencesReferences 1-50 are for the main text.
- Louie et al. [2021] S. G. Louie, Y.-H. Chan, F. H. da Jornada, Z. Li, and D. Y. Qiu, Discovering and understanding materials through computation, Nature Materials 20, 728 (2021).
- Merchant et al. [2023] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80 (2023).
- Li et al. [2020] J. Li, K. Lim, H. Yang, Z. Ren, S. Raghavan, P.-Y. Chen, T. Buonassisi, and X. Wang, Ai applications through the whole life cycle of material discovery, Matter 3, 393 (2020).
- Raabe et al. [2023] D. Raabe, J. R. Mianroodi, and J. Neugebauer, Accelerating the design of compositionally complex materials via physics-informed artificial intelligence, Nature Computational Science 3, 873 (2023).
- Nippa et al. [2023] D. F. Nippa, K. Atz, R. Hohler, A. T. Müller, A. Marx, C. Bartelmus, G. Wuitschik, I. Marzuoli, V. Jost, J. Wolfard, et al., Enabling late-stage drug diversification by high-throughput experimentation with geometric deep learning, Nature Chemistry , 1 (2023).
- Tropsha et al. [2023] A. Tropsha, O. Isayev, A. Varnek, G. Schneider, and A. Cherkasov, Integrating qsar modelling and deep learning in drug discovery: the emergence of deep qsar, Nature Reviews Drug Discovery , 1 (2023).
- Weiss et al. [2023] T. Weiss, E. Mayo Yanes, S. Chakraborty, L. Cosmo, A. M. Bronstein, and R. Gershoni-Poranne, Guided diffusion for inverse molecular design, Nature Computational Science 3, 873 (2023).
- Noé et al. [2020] F. Noé, A. Tkatchenko, K.-R. Müller, and C. Clementi, Machine learning for molecular simulation, Annual review of physical chemistry 71, 361 (2020).
- Wigh et al. [2022] D. S. Wigh, J. M. Goodman, and A. A. Lapkin, A review of molecular representation in the age of machine learning, Wiley Interdisciplinary Reviews: Computational Molecular Science 12, e1603 (2022).
- Choudhary et al. [2022] K. Choudhary, B. DeCost, C. Chen, A. Jain, F. Tavazza, R. Cohn, C. W. Park, A. Choudhary, A. Agrawal, S. J. Billinge, et al., Recent advances and applications of deep learning methods in materials science, npj Computational Materials 8, 59 (2022).
- Antoniuk et al. [2023] E. R. Antoniuk, G. Cheon, G. Wang, D. Bernstein, W. Cai, and E. J. Reed, Predicting the synthesizability of crystalline inorganic materials from the data of known material compositions, npj Computational Materials 9, 155 (2023).
- Noh et al. [2020] J. Noh, G. H. Gu, S. Kim, and Y. Jung, Machine-enabled inverse design of inorganic solid materials: promises and challenges, Chemical Science 11, 4871 (2020).
- Xiao et al. [2023] H. Xiao, R. Li, X. Shi, Y. Chen, L. Zhu, X. Chen, and L. Wang, An invertible, invariant crystal representation for inverse design of solid-state materials using generative deep learning, Nature Communications 14, 7027 (2023).
- Steinberger et al. [2019] D. Steinberger, H. Song, and S. Sandfeld, Machine learning-based classification of dislocation microstructures, Frontiers in Materials 6, 141 (2019).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- Zagorac et al. [2019] D. Zagorac, H. Müller, S. Ruehl, J. Zagorac, and S. Rehme, Recent developments in the Inorganic Crystal Structure Database: theoretical crystal structure data and related features, Journal of Applied Crystallography 52, 918 (2019).
- Curtarolo et al. [2012] S. Curtarolo, W. Setyawan, G. L. Hart, M. Jahnatek, R. V. Chepulskii, R. H. Taylor, S. Wang, J. Xue, K. Yang, O. Levy, M. J. Mehl, H. T. Stokes, D. O. Demchenko, and D. Morgan, Aflow: An automatic framework for high-throughput materials discovery, Computational Materials Science 58, 218 (2012).
- Vu et al. [2023] T.-S. Vu, M.-Q. Ha, D.-N. Nguyen, V.-C. Nguyen, Y. Abe, T. Tran, H. Tran, H. Kino, T. Miyake, K. Tsuda, et al., Towards understanding structure–property relations in materials with interpretable deep learning, npj Computational Materials 9, 215 (2023).
- Thirunavukarasu et al. [2023] A. J. Thirunavukarasu, D. S. J. Ting, K. Elangovan, L. Gutierrez, T. F. Tan, and D. S. W. Ting, Large language models in medicine, Nature medicine 29, 1930 (2023).
- Kasneci et al. [2023] E. Kasneci, K. Seßler, S. Küchemann, M. Bannert, D. Dementieva, F. Fischer, U. Gasser, G. Groh, S. Günnemann, E. Hüllermeier, et al., Chatgpt for good? on opportunities and challenges of large language models for education, Learning and individual differences 103, 102274 (2023).
- Romera-Paredes et al. [2023] B. Romera-Paredes, M. Barekatain, A. Novikov, M. Balog, M. P. Kumar, E. Dupont, F. J. Ruiz, J. S. Ellenberg, P. Wang, O. Fawzi, et al., Mathematical discoveries from program search with large language models, Nature 625, 468 (2023).
- Pereira et al. [2023] J. Pereira, R. Fidalgo, R. Lotufo, and R. Nogueira, Visconde: Multi-document qa with gpt-3 and neural reranking, in European Conference on Information Retrieval (Springer, 2023) pp. 534–543.
- Chen et al. [2024] D. Chen, S. Song, Q. Yu, Z. Li, W. Wang, F. Xiong, and B. Tang, Grimoire is all you need for enhancing large language models (2024), arXiv:2401.03385 .
- Xie et al. [2023] T. Xie, Y. Wa, W. Huang, Y. Zhou, Y. Liu, Q. Linghu, S. Wang, C. Kit, C. Grazian, and B. Hoex, Large language models as master key: Unlocking the secrets of materials science with gpt (2023), arXiv:2304.02213 .
- Zheng et al. [2023] Z. Zheng, O. Zhang, C. Borgs, J. T. Chayes, and O. M. Yaghi, Chatgpt chemistry assistant for text mining and the prediction of mof synthesis, Journal of the American Chemical Society 145, 18048 (2023).
- Min et al. [2023] B. Min, H. Ross, E. Sulem, A. P. B. Veyseh, T. H. Nguyen, O. Sainz, E. Agirre, I. Heintz, and D. Roth, Recent advances in natural language processing via large pre-trained language models: A survey, ACM Computing Surveys 56, 1 (2023).
- Devlin et al. [2018] J. Devlin, M.-W. Chang, K. Lee, and K. Toutanova, Bert: Pre-training of deep bidirectional transformers for language understanding (2018), arXiv:1810.04805 .
- Morgan and Jacobs [2020] D. Morgan and R. Jacobs, Opportunities and challenges for machine learning in materials science, Annual Review of Materials Research 50, 71 (2020).
- Deng et al. [2009] J. Deng, W. Dong, R. Socher, L.-J. Li, K. Li, and L. Fei-Fei, Imagenet: A large-scale hierarchical image database, in 2009 IEEE conference on computer vision and pattern recognition (Ieee, 2009) pp. 248–255.
- Fan et al. [2023] X. Fan, S. Chen, B. Steingrimsson, Q. Xiong, W. Li, and P. K. Liaw, Dataset for fracture and impact toughness of high-entropy alloys, Scientific Data 10, 37 (2023).
- Zhang et al. [2023a] Z. Zhang, H. Tang, and Z. Xu, Fatigue database of complex metallic alloys, Scientific Data 10, 447 (2023a).
- Gorsse et al. [2018] S. Gorsse, M. Nguyen, O. N. Senkov, and D. B. Miracle, Database on the mechanical properties of high entropy alloys and complex concentrated alloys, Data in brief 21, 2664 (2018).
- Li et al. [2023] X. Li, X. Hu, X. Chen, J. Fan, Z. Zhao, J. Wu, H. Wang, and Q. Dai, Spatial redundancy transformer for self-supervised fluorescence image denoising, Nature Computational Science 3, 1067 (2023).
- Zhou et al. [2023] G. Zhou, Z. Gao, Q. Ding, H. Zheng, H. Xu, Z. Wei, L. Zhang, and G. Ke, Uni-mol: a universal 3d molecular representation learning framework (2023).
- Varshneya and Mauro [2019] A. K. Varshneya and J. C. Mauro, Chapter 1 - introduction, in Fundamentals of Inorganic Glasses (Third Edition), edited by A. K. Varshneya and J. C. Mauro (Elsevier, 2019) third edition ed., pp. 1–18.
- Halim et al. [2021] Q. Halim, N. A. N. Mohamed, M. R. M. Rejab, W. N. W. A. Naim, and Q. Ma, Metallic glass properties, processing method and development perspective: a review, The International Journal of Advanced Manufacturing Technology 112, 1231 (2021).
- Liu et al. [2023a] G. Liu, S. Sohn, C. S. O’Hern, A. C. Gilbert, and J. Schroers, Effective subgrouping enhances machine learning prediction in complex materials science phenomena: Inoue’s subgrouping in discovering bulk metallic glasses, Acta Materialia 265, 119590 (2023a).
- Li et al. [2017] Y. Li, S. Zhao, Y. Liu, P. Gong, and J. Schroers, How many bulk metallic glasses are there?, ACS combinatorial science 19, 687 (2017).
- Zhou et al. [2022] Z. Zhou, Y. Shang, and Y. Yang, A critical review of the machine learning guided design of metallic glasses for superior glass-forming ability, Journal of Materials Informatics 2, 1 (2022).
- Ward et al. [2018] L. Ward, S. C. O’Keeffe, J. Stevick, G. R. Jelbert, M. Aykol, and C. Wolverton, A machine learning approach for engineering bulk metallic glass alloys, Acta Materialia 159, 102 (2018).
- Rodriguez et al. [2023] A. D. Rodriguez, K. R. Dearstyne, and J. Cleland-Huang, Prompts matter: Insights and strategies for prompt engineering in automated software traceability, in 2023 IEEE 31st International Requirements Engineering Conference Workshops (REW) (IEEE, 2023) pp. 455–464.
- Gupta et al. [2022] T. Gupta, M. Zaki, N. A. Krishnan, and Mausam, Matscibert: A materials domain language model for text mining and information extraction, npj Computational Materials 8, 102 (2022).
- Chakraborty et al. [2017] S. Chakraborty, R. Tomsett, R. Raghavendra, D. Harborne, M. Alzantot, F. Cerutti, M. Srivastava, A. Preece, S. Julier, R. M. Rao, et al., Interpretability of deep learning models: A survey of results, in 2017 IEEE smartworld, ubiquitous intelligence & computing, advanced & trusted computed, scalable computing & communications, cloud & big data computing, Internet of people and smart city innovation (IEEE, 2017) pp. 1–6.
- Ribeiro et al. [2016] M. T. Ribeiro, S. Singh, and C. Guestrin, ”Why Should I Trust You?”: Explaining the Predictions of Any Classifier (2016), arXiv:1602.04938 .
- Inoue [2000] A. Inoue, Stabilization of metallic supercooled liquid and bulk amorphous alloys, Acta materialia 48, 279 (2000).
- Ren et al. [2018] F. Ren, L. Ward, T. Williams, K. J. Laws, C. Wolverton, J. Hattrick-Simpers, and A. Mehta, Accelerated discovery of metallic glasses through iteration of machine learning and high-throughput experiments, Science advances 4, eaaq1566 (2018).
- Johnson et al. [2016] W. Johnson, J. Na, and M. Demetriou, Quantifying the origin of metallic glass formation, Nature communications 7, 10313 (2016).
- Wang [2009] W. Wang, Bulk metallic glasses with functional physical properties, Advanced Materials 21, 4524 (2009).
- Dawid and LeCun [2023] A. Dawid and Y. LeCun, Introduction to latent variable energy-based models: A path towards autonomous machine intelligence (2023), arXiv:2306.02572 .
- Hendrycks et al. [2021] D. Hendrycks, C. Burns, S. Basart, A. Zou, M. Mazeika, D. Song, and J. Steinhardt, Measuring massive multitask language understanding (2021), arXiv:2009.03300 .
- Joiner [2018] I. A. Joiner, Chapter 1 - artificial intelligence: Ai is nearby, in Emerging Library Technologies, Chandos Information Professional Series, edited by I. A. Joiner (Chandos Publishing, 2018) pp. 1–22.
- Niu et al. [2021] Z. Niu, G. Zhong, and H. Yu, A review on the attention mechanism of deep learning, Neurocomputing 452, 48 (2021).
- Wei et al. [2023] J. Wei, X. Wang, D. Schuurmans, M. Bosma, B. Ichter, F. Xia, E. Chi, Q. Le, and D. Zhou, Chain-of-thought prompting elicits reasoning in large language models (2023), arXiv:2201.11903 .
- Ray [2023] P. P. Ray, Chatgpt: A comprehensive review on background, applications, key challenges, bias, ethics, limitations and future scope, Internet of Things and Cyber-Physical Systems 3, 121 (2023).
- Helm et al. [2020] J. M. Helm, A. M. Swiergosz, H. S. Haeberle, J. M. Karnuta, J. L. Schaffer, V. E. Krebs, A. I. Spitzer, and P. N. Ramkumar, Machine learning and artificial intelligence: definitions, applications, and future directions, Current reviews in musculoskeletal medicine 13, 69 (2020).
- Almeida and Xexeo [2023] F. Almeida and G. Xexeo, Word embeddings: A survey (2023), arXiv:1901.09069 .
- Zheng and Casari [2018] A. Zheng and A. Casari, Feature Engineering for Machine Learning: Principles and Techniques for Data Scientists, 1st ed. (O’Reilly Media, Inc., 2018).
- Svozil et al. [1997] D. Svozil, V. Kvasnicka, and J. Pospichal, Introduction to multi-layer feed-forward neural networks, Chemometrics and intelligent laboratory systems 39, 43 (1997).
- Wang et al. [2020] Y. Wang, Q. Yao, J. T. Kwok, and L. M. Ni, Generalizing from a few examples: A survey on few-shot learning, ACM Comput. Surv. 53, 1 (2020).
- Natekin and Knoll [2013] A. Natekin and A. Knoll, Gradient boosting machines, a tutorial, Frontiers in neurorobotics 7, 21 (2013).
- Yang and Shami [2020] L. Yang and A. Shami, On hyperparameter optimization of machine learning algorithms: Theory and practice, Neurocomputing 415, 295 (2020).
- Dong et al. [2023] Q. Dong, L. Li, D. Dai, C. Zheng, Z. Wu, B. Chang, X. Sun, J. Xu, L. Li, and Z. Sui, A survey on in-context learning (2023), arXiv:2301.00234 .
- Gou et al. [2021] J. Gou, B. Yu, S. J. Maybank, and D. Tao, Knowledge distillation: A survey, International Journal of Computer Vision 129, 1789 (2021).
- Brown et al. [2020] T. B. Brown, B. Mann, N. Ryder, M. Subbiah, J. Kaplan, P. Dhariwal, A. Neelakantan, P. Shyam, G. Sastry, A. Askell, S. Agarwal, A. Herbert-Voss, G. Krueger, T. Henighan, R. Child, A. Ramesh, D. M. Ziegler, J. Wu, C. Winter, C. Hesse, M. Chen, E. Sigler, M. Litwin, S. Gray, B. Chess, J. Clark, C. Berner, S. McCandlish, A. Radford, I. Sutskever, and D. Amodei, Language models are few-shot learners (2020), arXiv:2005.14165 .
- Zhang et al. [2023b] Y. Zhang, Y. Li, L. Cui, D. Cai, L. Liu, T. Fu, X. Huang, E. Zhao, Y. Zhang, Y. Chen, L. Wang, A. T. Luu, W. Bi, F. Shi, and S. Shi, Siren’s song in the ai ocean: A survey on hallucination in large language models (2023b), arXiv:2309.01219 .
- Menard [2002] S. Menard, Applied logistic regression analysis, 106 (Sage, 2002).
- Han et al. [2021] X. Han, Z. Zhang, N. Ding, Y. Gu, X. Liu, Y. Huo, J. Qiu, Y. Yao, A. Zhang, L. Zhang, W. Han, M. Huang, Q. Jin, Y. Lan, Y. Liu, Z. Liu, Z. Lu, X. Qiu, R. Song, J. Tang, J.-R. Wen, J. Yuan, W. X. Zhao, and J. Zhu, Pre-trained models: Past, present and future (2021), arXiv:2106.07139 .
- Wang et al. [2023] Y. Wang, Y. Kordi, S. Mishra, A. Liu, N. A. Smith, D. Khashabi, and H. Hajishirzi, Self-instruct: Aligning language models with self-generated instructions (2023), arXiv:2212.10560 .
- Antonelli et al. [2019] L. Antonelli, M. R. Guarracino, L. Maddalena, and M. Sangiovanni, Integrating imaging and omics data: a review, Biomedical Signal Processing and Control 52, 264 (2019).
- Suthaharan [2016] S. Suthaharan, Support vector machine, in Machine Learning Models and Algorithms for Big Data Classification: Thinking with Examples for Effective Learning (Springer US, Boston, MA, 2016) pp. 207–235.
- Kudo and Richardson [2018] T. Kudo and J. Richardson, Sentencepiece: A simple and language independent subword tokenizer and detokenizer for neural text processing (2018).
- Pan and Yang [2009] S. J. Pan and Q. Yang, A survey on transfer learning, IEEE Transactions on knowledge and data engineering 22, 1345 (2009).
- Tal [2011] J. Tal, Chapter 14 - sample size, in Strategy and Statistics in Clinical Trials, edited by J. Tal (Academic Press, Boston, 2011) pp. 229–244.
- Liu et al. [2023b] P. Liu, W. Yuan, J. Fu, Z. Jiang, H. Hayashi, and G. Neubig, Pre-train, prompt, and predict: A systematic survey of prompting methods in natural language processing, ACM Computing Surveys 55, 1 (2023b).
- Yao et al. [2023] S. Yao, J. Zhao, D. Yu, N. Du, I. Shafran, K. Narasimhan, and Y. Cao, React: Synergizing reasoning and acting in language models (2023), arXiv:2210.03629 .
- Lewis et al. [2021] P. Lewis, E. Perez, A. Piktus, F. Petroni, V. Karpukhin, N. Goyal, H. Kttler, M. Lewis, W. tau Yih, T. Rocktaschel, S. Riedel, and D. Kiela, Retrieval-augmented generation for knowledge-intensive nlp tasks (2021), arXiv:2005.11401 .
- Vaswani et al. [2023] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, L. Kaiser, and I. Polosukhin, Attention is all you need (2023), arXiv:1706.03762 .
- Beltagy et al. [2020] I. Beltagy, M. E. Peters, and A. Cohan, Longformer: The long-document transformer (2020), arXiv:2004.05150 .
- cohere docs [2024a] cohere docs, Top-k and top-p, https://docs.cohere.com/docs/controlling-generation-with-top-k-top-p (accessed on January 29, 2024a).
- cohere docs [2024b] cohere docs, Temperature, https://docs.cohere.com/docs/temperature (accessed on January 29, 2024b).
Acknowledgments
This work is supported by Research Grants Council, Hong Kong SAR through the Collaborative Research Fund (C1005-19G) and General Research Fund (17210723). T.W. acknowledges additional support by The University of Hong Kong (HKU) via seed funds (2201100392, 2309100163). S.P. acknowledges additional support by HKU via seed fund (2309100201). We acknowledge helpful discussions with Prof. Yi Ma at HKU and Simon Zhai at UC Berkeley.
Competing interests
The authors declare no competing interests.
Supplementary Information
Supplementary Notes
Supplementary Figures S1-15
Supplementary Tables S1-2
References (51-80)
Supplementary Information for
A Prompt-Engineered Large Language Model, Deep Learning Workflow
for Materials Classification
Siyu Liu et al.
1 Supplementary Notes
1.1 Glossary
Artificial Intelligence: artificial intelligence (AI) refers to the development of computer systems capable of performing tasks that typically require human intelligence, including visual perception, decision-making, and language comprehension [51].
Attention: a mechanism enabling models to focus on specific parts of the input sequence during processing, allowing them to weigh the significance of different words when generating output. This attention mechanism assists in capturing long-range dependencies and enhancing performance in various natural language processing (NLP) tasks [52].
BERT: Bidirectional Encoder Representations from Transformers, a widely used pre-trained language model [27].
Chain-of-thoughts: a prompt engineering method that enhances complex reasoning capabilities by adding intermediate reasoning steps to the text input for the model [53].
ChatGPT: Chat Generative Pre-trained Transformer, developed by OpenAI, is a chatbot and a sibling model to InstructGPT, trained to follow instructions in a prompt and deliver detailed responses [54].
Cross Entropy Loss: a performance indicator measuring classification models, also known as log loss, with lower values indicating better predictive performance.
Deep Learning: deep learning (DL) is an AI technology emulating the neural network structure of the human brain, learning and training via multi-level neurons for complex data analysis and processing [55].
Embedding: an NLP technique that maps a high-dimensional space, with a dimensionality equal to the total number of words, to a much lower-dimensional continuous vector space, assigning each word or phrase a vector in the real number domain [56].
Feature Engineering: the process of selecting, transforming, and creating new features from raw data to improve machine leanring (ML) model performance. It involves identification of relevant and informative input variables and their preparation for accurate predictions or classifications [57].
Feed Forward: refers to a deep learning framework, also known as a multi-layer perceptron (MLP), that defines a mapping , and learns the value of parameter to achieve the best function approximation [58].
Few-shot Learning: a ML approach where models are trained to make accurate predictions with limited examples per class [59].
Gemini: Google’s multimodal large language model (LLM), the first to outperform human experts on MMLU (Massive Multitask Language Understanding), a popular method for testing AI models’ knowledge and problem-solving abilities.
Gradient Boosting Decision Tree (GBDT): also known as Gradient Boosting Machines (GBM), is a popular ensemble learning method combining decision trees with gradient boosting. GBDT continuously fits new models for more accurate estimates of response variables. The principle behind this algorithm is to construct new base learners that maximizes correlation with the negative gradient of the ensemble’s loss function [60].
Hyperparameters: external configuration variables input in advance to manage the training of ML models [61].
In-context Learning: refers to enhancing LLM performance using a few examples provided in the input context [62].
ImageNet: an image database organized according to the WordNet hierarchy (currently only nouns), with each hierarchy node represented by hundreds of thousands of images [29].
Knowledge Distillation: refers to extracting knowledge from larger deep neural networks into smaller networks [63].
Large Language Model: large language model (LLM) is a kind of DL model trained on large amounts of text data to learn statistical patterns of natural language [64].
LLM Hallucination: occurs when LLMs generate content deviating from user input, contradicting previously generated context, or inconsistent with established world knowledge [65].
Logistic Regression: a supervised learning algorithm utilizing logistic functions to estimate label probabilities [66].
Machine Learning: machine learning (ML) demonstrates the experiential ‘learning’ associated with human intelligence, along with the ability to enhance its analyses through using computational algorithms [55].
Multi-head Attention: represents multiple attention modules within a single attention layer of the model, allowing for different focus on various parts of a sequence.
Pre-trained Model: a model trained on a large corpus of data and can be fine-tuned to solve various tasks [67].
Prompt Engineering: a method to generate textual data from LLMs by embedding task descriptions in the input, effectively conveying specific parameters to the model as part of a problem statement [54].
Self-instruct Learning: a method for improving the instruction-following capabilities of pre-trained language models by bootstrapping off their own generations [68].
Supervised Classification: a ML task where the goal is to categorize input data into predefined classes or categories based on labeled training examples [69].
Support Vector Machine: a supervised learning algorithm for classification and regression tasks that identifies the optimal boundary (hyperplane) separating data points of different classes [70].
Tokenization: the process of converting text into smaller structural markers, called tokens. The tool used to handle this process is known as a tokenizer [71].
Transfer Learning: a ML technique employing a pre-trained model as the starting point for a new related task, instead of training a model from scratch. It allows for different domains, tasks, and distributions in training and testing [72].
1.2 Workflow details
1.2.1 Data processing of metallic glasses dataset
As mentioned in the “Methods” section, our initial dataset consists of 5,577 samples. To train different BERT models, it is necessary to divide it into a training and a test set. We employed a stratified sampling method for this purpose, ensuring a representative and balanced representation of the different subgroups [73], which enhances the quality and reliability of sample. Based on the overall 80:20 split ratio for the training and test sets, we have performed non-repetitive stratified sampling on three alloy data categories. In the end, we obtained a training set with 4,460 samples and a test set with 1,117 samples, maintaining the 80:20 ratio for each category.
It should be noted that this section aims to establish the relationship between alloy composition and labels. The textual data generated through Gemini-pro is used to train different BERT classification models. The data obtained in this section is stored in the “original data” folder, with the data format shown in Table S1.
1.2.2 Prompt design and textual data generation
Upon acquiring the alloy composition data, we proceeded with prompt design. Prompt engineering is a prominent direction in the field of LLMs and is considered a method that significantly improve data generation effectiveness for LLMs [74]. Techniques such as few-shot learning [53], CoT [53], synergizing reasoning and acting (ReAct) [75], and retrieval-augmented generation (RAG) [76] are deemed effective prompt engineering methods. Due to the additional knowledge sources required for the latter two methods, we tested direct inquiry, few-shot, and CoT, and compared them with our MetalPrompt. Figs. S8, S9, and S10 are templates for three benchmark methods, while Fig. S4 shows the comparison results. Our “MetalPrompt” achieved the best performance.
For the prompt inquiry setting with the LLM, we used Top-K = 1 and Temperature = 0 as default parameters. As shown in Fig. S11 and Fig. S12, a smaller Top-K and lower temperature indicate more reliable model output. Top-K = 1 implies that the model will perform greedy decoding and select the most probable value. Unless specifically explained, the parameters of the language model mentioned below are consistent with these settings.
Using the above parameter settings and “MetalPrompt” as the input template, we generated textual data for 5,577 alloy compositions. The replacement of alloy elements in the template was assisted by the langchain library. Simultaneously, based on the alloy composition category relationship of the original training and test dataset, we labeled the generated textual data with the same category. This dataset is used for training the classification model below, with the partitioning of training and testing sets consistent with the original data.
1.2.3 MgBERT training and evaluation
BERT model and attention mechanism: BERT, introduced by Google researchers in 2018 [27], is a pre-training language representation method that significantly impacted the field of NLP. Traditional language models analyzed text data unidirectionally, either from left to right or right to left, limiting their understanding of language context. BERT, employing the Transformer architecture, processes each word concerning all other words in a sentence, rather than sequentially. Owing to its state-of-the-art results in numerous tasks, BERT was selected as the foundation for our classification model.
Fig. 4c in the main text shows the basic BERT architecture, comprising data pre-processing, input encoding, model training, post-processing, and output results. During data processing, tokenizers divide sentences or words into individual tokens, such as marking “for example” as “for” and “example”. Then, as demonstrated in Fig. S13, the encoding method transforms pre-processed text data into the input representation of the model. For instance, “for” is converted to a vector . Model training is crucial to BERT, as it uses multi-head attention to extract essential features from input embeddings. Attention mechanisms enhance model performance in NLP and other sequence data processing tasks by focusing on the most relevant parts of the input sequence. Models can learn to assign various attention weights to input information from different positions, concentrating on crucial aspects when processing input sequences. This mechanism allows the model to capture long-distance dependencies and important patterns in sequences more effectively, thus improving its performance in processing sequence data [77]. Fig. S14 shows a schematic diagram of the operation of the attention mechanism. Multi-head attention, compared to a single module, employs multiple attention modules in the same calculation, extracting different feature information. The calculation formula for attention score is as follows:
Where .
In Equation S1, refers to the queries and dimension keys. , , and refer to the input query, key, and value, respectively. refers to the softmax function.
Pre-trained models selection: We compared three pre-trained models based on BERT, including the basic BERT model [27], Longformer supporting longer input text lengths [78], and MatSciBERT trained on materials science texts [42].
All models employ default parameter settings; however, to expedite calculations, the maximum input length for Longformer is limited to 1200 tokens, while other models have a default maximum length of 512. To ensure reproducibility, the random seed is fixed at 42. All models are trained using cross-entropy loss and optimized with the Adam optimizer for parameters. All training was conducted on a 32G Nvidia V100. Table S2 presents the basic hyperparameters for model training. Due to its larger embedding layer requiring more graphics memory, Longformer has a slightly smaller batch size.
MgBERT embedding and training: Upon the training of MgBERT, we found that the file with the longest tokens is ’Fe68.3C6.9Si2.5B6.7P8.8Cr2.2Mo2.5Al2.1.txt’ with 859 tokens, and 300 files have more than 512 tokens. Our basic model, MatSciBERT, only supports inputs with a maximum token length of 512. Considering the potential inconsistency in text input length for different materials and scalability requirements, we designed a variable-length MgBERT position embedding layer that does not necessitate complete retraining. The following Algorithm 1 is a pseudocode representation of its principle.
For the comparison of three embedding methods, we ran 16 epochs with a learning rate of 3e-5 and a batch size of 26. In Figs. 4g, h of the main text, only the first 15 epochs are displayed due to the loss output retaining the first three decimal places. In the 16th epoch, some models exhibit training set loss 0.001. The complete training information can be found in the “train log” folder.
For the final MgBERT model, we continued training with the highest classification accuracy model weight of Method 3 at a learning rate of 2e-5 and compared it to the case with 1e-5. The highest accuracy of 88.5% was achieved when the learning rate was 2e-5 and the epoch was 4. We utilized the model weights trained under these parameters as the final MgBERT model.
Model evaluation: During the model evaluation stage, we compared some ML models, with their concepts outlined in the “Glossary” section. Model implementation is based on the “scikit-learn” software and employs its default hyperparameters for training. The random seed is fixed at 42 for reproducibility. The training and test set rely on stratified sampling, as described in the “Methods” section of the main text.
1.2.4 Model interpretability
We used the LIME model [44] for model interpretation. LIME utilizes a trained local surrogate model to explain individual predictions. For a black-box model of interest, LIME samples around the instance of interest and collects perturbed instances along with their black-box model predictions. A local interpretable model (e.g. linear regression, decision tree) is then trained on this new dataset to approximate the behavior of the black-box model locally. Fig. S15 from [44] shows its working principle, wherein the dashed line is a trained linear model used to explain the black-box model with complex boundaries in the red and blue part of the graph. The hyperparameters we set here is to perform 1,000 samples and filter the top 50 contributing features. Considering the maximum token length does not exceed 900, this choice is reasonable under the limitation of computational rate, as the training time increases exponentially with the increase of sampling times.
Additionally, we visualized the attention score of the [CLS] (classify) token in the last layer. Different attention scores represent the degree of importance that a specific token places on other tokens, with larger attention scores indicating higher importance. The classifier processes the value of the last layer’s [CLS] token, compressing this 768-dimensional vector into three dimensions. Each dimension represents the probability of different categories. Therefore, calculating the focus on the previous layer assists in interpreting the model.
2 Supplementary figures
3 Supplementary Tables
glass forming category |
composition |
Ribbon |
Ag20Al25La55 |
BMG |
Cu55Zr42.5Ga2.5 |
NR |
Ag6Ce8Cu86 |
model name |
hyperparameters |
BERT |
Epoch: 8; Learning rate: 3e-5; Batch size: 26 |
longformer |
Epoch: 8; Learning rate: 3e-5; Batch size: 12 |
MatSciBERT |
Epoch: 8; Learning rate: 3e-5; Batch size: 26 |